# OAC: Output-adaptive Calibration for Accurate Post-training Quantization

### v1

@@ 6 Experiments and Results
This section provides the key experimental results in addition to a summary of the settings used in our experiments. We refer to the appendix for further discussion of the findings and settings.
@@ 6.1 Experimental Settings
We investigate various language models with different sizes including OPT [42], LLaMa [36], and LLaMa 2 [37] families. The calibration set comprises 128 sequences of 2048 tokens. To evaluate the performance of the quantized models on language modeling tasks, we report their perplexity on C4 [31] and WikiText2 [27]. Also, Language Model Evaluation Harness (LMEH) [12] is utilized for evaluating the reasoning abilities of the quantized models. We report the zero-shot accuracy on WinoGrande [32], PiQA [35], HellaSwag [41], ARC-easy, and ARC-challenge [4] in addition to the five-shot exact match on GSM8K [5] datasets. See Appendix F for more details on the datasets.
In our experimental results, we compare our method with the latest state-of-the-art PTQ methods that are developed for LLMs including Round to Nearest (RTN) [7], OPTQ [11], QuIP [3], and SpQR [8] on 2- and 3- bit quantization in addition to OmniQuant [34] as an efficient QAT method. SqueezeLLM [19] is also included in the 3-bit quantization experiments. Moreover, BiLLM [15] is selected as the most recent baseline for binary PTQ. Detailed information on the experimental settings such as the quantization configurations are provided in the Appendix G.
@@ 6.2 Results
Our experimental results reveal the importance of output-adaptive calibration in low-precision quantization scenarios. Our proposed approach performs better than other methods, especially, in more complicated scenarios where the average bit width or model size is relatively smaller.
@@ C.2 Hessian Regularization
When employing the Hessian inverse for the calibration in PTQ, a regularization term is added to the diagonal elements of the Hessian matrix. This additional term serves to mitigate numerical instability, which is common in computing the inverse of matrices [11, 3, 8, 15].
@@ C.3 Hessian Reduction Over Calibration Samples
$\widehat{\mathbf{H}}_{\mathrm{OAC}}=\sum_{i=1}^{N}{\mathbf{G}_{i}^{\top}\mathbf{G}_{i}}$ (19)
@@ Appendix F Datasets
We create the calibration sets used by the studied PTQ methods by randomly selecting 128 samples of length 2048 from C4 [31], RedPajama [6], or WikiText2 [27] datasets. The C4 samples are mostly used for the calibration of OPT models. The LLaMa 1 and LLaMa 2 models are calibrated mostly based on the RedPajama samples. However, some PTQ baselines do not support using RadPajama in their official implementation for quantizing the LLaMa models. In those cases, C4 samples are used instead of the RadPajama samples for the calibration of such methods. Also, the WikiText2 samples are used in the OmniQuant experiments for calibrating the OPT and, LLaMa 1, and LLaMa 2 families. Refer to Section G for the detailed implementation configurations of each studied method.
@@
The WikiText2 [27] and PTB [26] perplexities are reported on their standard test sets. The test dataset for reporting the C4 perplexity is created by randomly selecting 1024 data samples from the C4 validation set. We selected WinoGrande [32], PiQA [35], HellaSwag [41], GSM8K [5], ARC-easy, and ARC-challenge [4] datasets from Language Model Evaluation Harness (LMEH) [12] to evaluate the reasoning abilities of the quantized models. We emphasize that GSM8K is excluded from the OPT evaluations due to the poor performance of uncompressed OPT models on this challenging task while we keep GSM8K in the evaluation of the quantized LLaMa models.
@@ Appendix G Quantization Configurations
We use up to 8$\times$×NVIDIA V100-32G GPUs for our experiments. Also, the Stochastic Gradient Descent (SGD) optimizer implemented by PyTorch [29] is used in our experiments to compute the gradients of the linear layers. We implemented our experiments using the Hugging Face transformers [38] framework. Here, the exact experimental configurations are reported for the baseline PTQ methods that are studied in our experiments.
@@
We implemented this method based on the SpQR official implementation [^3] See https://github.com/Vahe1994/SpQR. for RTN [7]. However, we integrated group quantization in our version of RTN to significantly improve the accuracy. RTN is applied in the 2- and 3-bit PTQ experiments and the quantization group size is set to 128.
@@
This method is implemented using the SpQR official implementation. We set the quantization group size to 128 to apply OPTQ [11] in the 2- and 3-bit PTQ experiments. The calibration samples are drawn from C4 and RedPajama to quantize the OPT and LLaMa models, respectively. We also tune the Hessian regularization factor, $\alpha$ for this method.
@@
This method is implemented using its Offical implementation [^4] See https://github.com/OpenGVLab/OmniQuant. in 2- and 3-bit weight-only quantization experiments. As proposed in the paper [34], 128 samples from WikiText2 are randomly selected to compose the calibration set and the quantization group size is set to 128. We reproduce the quantized OPT models using this method following the hyperparameters recommended in their paper. To employ OmniQuant for the quantization of the LLaMa 1 and LLaMa 2 models, we download and evaluate the quantized checkpoints provided in the official public repository.
@@
We use the SpQR official implementation to apply SpQR in the binary, 2-bit, and 3-bit PTQ experiments. Table 8 shows the hyperparameters that are used for running the SpQR [8] experiments. We also tune the Hessian regularization factor, $\alpha$ for this method.
@@
This method is implemented using its official implementation [^5] See https://github.com/Cornell-RelaxML/QuIP.. We tune $\alpha$ and draw the calibration samples from C4 [3] in the 2- and 3-bit PTQ experiments.
@@
We use the quantized checkpoints provided by the SqueezeLLM [19] official implementation [^6] See https://github.com/SqueezeAILab/SqueezeLLM. to reproduce this method in the 3-bit PTQ experiments.
@@
This method [15] is applied for binary PTQ of the LLaMa 1 and LLaMA 2 models using its official implementation [^7] See https://github.com/Aaronhuang-778/BiLLM.. We tune the Hessian regularization factor, $\alpha$ for this method. Also, the calibration set is selected from C4.

### v2

@@ 6 Experiments and Results
This section provides the key experimental results in addition to a summary of the settings used in our experiments. We refer to the appendix for further discussion of the findings, hyper-parameters, and settings.
@@ 6.1 Experimental Settings
We investigate various language models with different sizes including OPT (Zhang et al., 2022), LLaMa (Touvron et al., 2023a), and LLaMa 2 (Touvron et al., 2023b) families. The calibration set comprises 128 sequences of 2048 tokens. To evaluate the performance of the quantized models on language modeling tasks, we report their perplexity on C4 (Raffel et al., 2020) and WikiText2 (Merity et al., 2017). Also, Language Model Evaluation Harness (LMEH) (Gao et al., 2023) is utilized for evaluating the reasoning abilities of the quantized models. We report the zero-shot accuracy on WinoGrande (Sakaguchi et al., 2021), PiQA (Tata and Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy, and ARC-challenge (Clark et al., 2018) in addition to the five-shot exact match on GSM8K (Cobbe et al., 2021) datasets. See Appendix F for more details on the datasets.
In our experimental results, we compare our method with the latest state-of-the-art PTQ methods that are developed for LLMs including Round to Nearest (RTN) (Dettmers et al., 2022), OPTQ (Frantar et al., 2023), QuIP (Chee et al., 2023), and SpQR (Dettmers et al., 2024) on 2- and 3- bit quantization in addition to OmniQuant (Shao et al., 2024) as an efficient QAT method. SqueezeLLM (Kim et al., 2024) is also included in the 3-bit quantization experiments. Moreover, BiLLM (Huang et al., 2024) is selected as the most recent baseline for binary PTQ. Detailed information on the experimental settings such as the quantization configurations are provided in the Appendix G.
@@ 6.2 Results
The experimental results reveal the importance of output-adaptive calibration in low-precision quantization scenarios. Our proposed approach, OAC, achieves a better accuracy compared to the other methods, especially, in more complicated scenarios where the average bit width or model size is relatively smaller.
@@ C.2 Hessian Regularization
When employing the Hessian inverse for the calibration in PTQ, a regularization term is added to the diagonal elements of the Hessian matrix. This additional term serves to mitigate numerical instability, which is common in computing the inverse of matrices [Frantar et al., 2023, Chee et al., 2023, Dettmers et al., 2024, Huang et al., 2024].
@@ C.3 Hessian Reduction Over Calibration Samples
$\widehat{\mathbf{H}}_{\mathrm{OAC}}=\sum_{i=1}^{N}{\mathbf{G}^{\top}[i]\mathbf{G}[i]}$ (19)
@@ Appendix F Datasets
We create the calibration sets used by the studied PTQ methods by randomly selecting 128 samples of length 2048 from C4 [Raffel et al., 2020], RedPajama [Computer, 2023], or WikiText2 [Merity et al., 2017] datasets. The C4 samples are mostly used for the calibration of OPT models. The LLaMa 1 and LLaMa 2 models are calibrated mostly based on the RedPajama samples. However, some PTQ baselines do not support using RadPajama in their official implementation for quantizing the LLaMa models. In those cases, C4 samples are used instead of the RadPajama samples for the calibration of such methods. Also, the WikiText2 samples are used in the OmniQuant experiments for calibrating the OPT and, LLaMa 1, and LLaMa 2 families. Refer to Section G for the detailed implementation configurations of each studied method.
@@
The WikiText2 [Merity et al., 2017] and PTB [Marcus et al., 1993] perplexities are reported on their standard test sets. The test dataset for reporting the C4 perplexity is created by randomly selecting 1024 data samples from the C4 validation set. We selected WinoGrande [Sakaguchi et al., 2021], PiQA [Tata and Patel, 2003], HellaSwag [Zellers et al., 2019], GSM8K [Cobbe et al., 2021], ARC-easy, and ARC-challenge [Clark et al., 2018] datasets from Language Model Evaluation Harness (LMEH) [Gao et al., 2023] to evaluate the reasoning abilities of the quantized models. We emphasize that GSM8K is excluded from the OPT evaluations due to the poor performance of uncompressed OPT models on this challenging task while we keep GSM8K in the evaluation of the quantized LLaMa models.
@@ Appendix G Quantization Configurations
We use up to 8$\times$×NVIDIA V100-32G GPUs for our experiments. Also, the Stochastic Gradient Descent (SGD) optimizer implemented by PyTorch [Paszke et al., 2019] is used in our experiments to compute the gradients of the linear layers. We implemented our experiments using the Hugging Face transformers [Wolf et al., 2020] framework. Here, the exact experimental configurations are reported for the baseline PTQ methods that are studied in our experiments.
@@
We implemented this method based on the SpQR official implementation [^3] See https://github.com/Vahe1994/SpQR. for RTN [Dettmers et al., 2022]. However, we integrated group quantization in our version of RTN to significantly improve the accuracy. RTN is applied in the 2- and 3-bit PTQ experiments and the quantization group size is set to 128.
@@
This method is implemented using the SpQR official implementation. We set the quantization group size to 128 to apply OPTQ [Frantar et al., 2023] in the 2- and 3-bit PTQ experiments. The calibration samples are drawn from C4 and RedPajama to quantize the OPT and LLaMa models, respectively. We also tune the Hessian regularization factor, $\alpha$ for this method.
@@
This method is implemented using its Offical implementation [^4] See https://github.com/OpenGVLab/OmniQuant. in 2- and 3-bit weight-only quantization experiments. As proposed in the paper [Shao et al., 2024], 128 samples from WikiText2 are randomly selected to compose the calibration set and the quantization group size is set to 128. We reproduce the quantized OPT models using this method following the hyperparameters recommended in their paper. To employ OmniQuant for the quantization of the LLaMa 1 and LLaMa 2 models, we download and evaluate the quantized checkpoints provided in the official public repository.
@@
We use the SpQR official implementation to apply SpQR in the binary, 2-bit, and 3-bit PTQ experiments. Table 8 shows the hyperparameters that are used for running the SpQR [Dettmers et al., 2024] experiments. We also tune the Hessian regularization factor, $\alpha$ for this method.
@@
This method is implemented using its official implementation [^5] See https://github.com/Cornell-RelaxML/QuIP.. We tune $\alpha$ and draw the calibration samples from C4 [Chee et al., 2023] in the 2- and 3-bit PTQ experiments.
@@
We use the quantized checkpoints provided by the SqueezeLLM [Kim et al., 2024] official implementation [^6] See https://github.com/SqueezeAILab/SqueezeLLM. to reproduce this method in the 3-bit PTQ experiments.
@@
This method [Huang et al., 2024] is applied for binary PTQ of the LLaMa 1 and LLaMA 2 models using its official implementation [^7] See https://github.com/Aaronhuang-778/BiLLM.. We tune the Hessian regularization factor, $\alpha$ for this method. Also, the calibration set is selected from C4.
@@ Appendix H Detailed Experimental Results
Appendix I Choice of the Hessian-based Calibration Method
[…]
In this section, we perform an ablation study to investigate if OAC consistently improves the accuracy if integrated with all of the studied Hessian-based calibration methods.
Table 14 shows the perplexity and LMEH average score of LLaMa-7B and LLaMa2-7B when quantized using the vanilla Hessian-based Calibration methods compared to integrating these methods to OAC.
Based on our results, presented in Table 14, integrating each Hessian-based calibration method into our proposed method, OAC, significantly improves the performance. These results also show the strength of our proposed Hessian computation approach.
Model Method Avg Bits C4 $\downarrow$ WikiText2 $\downarrow$ LMEH $\uparrow$
LLaMa-7B OPTQ 2.25 76.30 103.80 33.10
[…]
QuIP 2 19.2 17.69 39.38
[…]
SpQR 2.09 12.56 10.59 43.73
[…]
BiLLM 1.09 27.82 30.73 35.88
[…]
LLaMa2-7B OPTQ 2.25 147.46 201.8 30.46
[…]
QuIP 2 51.22 67.92 32.06
[…]
SpQR 2.09 13.22 11.09 42.90
[…]
BiLLM 1.08 28.00 25.59 35.61
[…]
Table 14: Comparison of the perplexity and reasoning score for the quantized LLaMa-7B and LLaMa2-7B models when the different Hessian-based calibration methods are integrated into OAC.